OpenStudy (anonymous):

(g + 4) – 3g = 1 + g?

OpenStudy (anonymous):

combine likes terms. g - 3g +4 = 1 + g -2g = -3 +g -3g = -3 divide both sides by -3 to get?

OpenStudy (anonymous):

thats true

OpenStudy (anonymous):

g = 1 g = 4 no solution identity

OpenStudy (anonymous):

?

OpenStudy (anonymous):

g=1

OpenStudy (anonymous):

cmon man really